# Nonreciprocal entanglement in cavity magnomechanics exploiting chiral cavity-magnon coupling

### v3

@@ VI Conclusion
We present a scheme for achieving nonreciprocal microwave-magnon and -phonon bipartite entanglements and photon-magnon-phonon tripartite entanglement in a CMM system based on the chiral cavity-magnon coupling. This is realized in a torus-shaped microwave cavity by placing the YIG sphere on a special line in the cavity to break the time-reversal symmetry of the two degenerate counter-propagating cavity modes. We analyze the effects of some major experimental imperfections, such as the coupling between the two degenerate cavity modes and the simultaneous couplings of the magnon mode to the two degenerate modes, and find that the nonreciprocity of the entanglements exhibits strong robustness towards these imperfections. The work may find applications of the CMM systems in noise-tolerant quantum processing, channel multiplexing quantum teleportation, and chiral magnonic quantum networks.
@@ ACKNOWLEDGEMENTS
We thank H. Jing for valuable feedback on the manuscript, and Z.-Q. Wang and R.-C. Shen for helpful discussions. This work has been supported by the National Key Research and Development Program of China (Grant No. 2022YFA1405200) and National Natural Science Foundation of China (Grant No. 92265202).
APPENDIX
Including the effect of the backscattering induced coupling between the two degenerate counter-propagating cavity modes, the Hamiltonian of the system becomes
$\begin{split}H^{\prime}/\hbar=&\sum_{j=\circlearrowright,\circlearrowleft}\omega_{a}a_{j}^{\dagger}a_{j}+\omega_{m}m^{\dagger}m+\frac{\omega_{b}}{2}(q^{2}+p^{2})+g_{m}m^{\dagger}mq\\ &+\sum_{j=\circlearrowright,\circlearrowleft}g_{j}(a_{j}^{\dagger}m+a_{j}m^{\dagger})+J(a_{\circlearrowright}a_{\circlearrowleft}^{\dagger}+a_{\circlearrowright}^{\dagger}a_{\circlearrowleft})+H_{d}/\hbar,\end{split}$ (A1)
where $J$ is the coupling strength between the two degenerate modes.
For the case where the microwave drive is applied to the CW circulating cavity mode, the linearized QLEs describing the quadrature fluctuations of the system can be written in the matrix form of
$\dot{u}(t)=A^{\prime}u(t)+n(t),$ (A2)
$A^{\prime}=\begin{pmatrix}-\kappa_{a}&\Delta_{\circlearrowright}&0&J&0&g_{\circlearrowright}&0&0\\ -\Delta_{\circlearrowright}&-\kappa_{a}&-J&0&-g_{\circlearrowright}&0&0&0\\ 0&J&-\kappa_{a}&\Delta_{\circlearrowleft}&0&g_{\circlearrowleft}&0&0\\ -J&0&-\Delta_{\circlearrowleft}&-\kappa_{a}&-g_{\circlearrowleft}&0&0&0\\ 0&g_{\circlearrowright}&0&g_{\circlearrowleft}&-\kappa_{m}&\tilde{\Delta}_{m}&\mathrm{Im}\,G_{m}&0\\ -g_{\circlearrowright}&0&-g_{\circlearrowleft}&0&-\tilde{\Delta}_{m}&-\kappa_{m}&-\mathrm{Re}\,G_{m}&0\\ 0&0&0&0&0&0&0&\omega_{b}\\ 0&0&0&0&-\mathrm{Re}\,G_{m}&-\mathrm{Im}\,G_{m}&-\omega_{b}&-\gamma_{b}\end{pmatrix},$ (A3)
$\langle m\rangle=\frac{E[(g_{\circlearrowright}\Delta_{a}-g_{\circlearrowleft}J)-ig_{\circlearrowright}\kappa_{a}]}{(\kappa_{m}J^{2}+\kappa_{a}\epsilon_{1}-\Delta_{a}\epsilon_{2})+i(\tilde{\Delta}_{m}J^{2}-2Jg_{\circlearrowright}g_{\circlearrowleft}+\Delta_{a}\epsilon_{1}+\kappa_{a}\epsilon_{2})},$ (A4)
For the case where the CCW circulating cavity mode is driven, the drift matrix takes the same form as in Eq. (A3), while the steady-state average $\langle m\rangle$ takes a slightly different form of
$\langle m\rangle=\frac{E[(g_{\circlearrowleft}\Delta_{a}-g_{\circlearrowright}J)-ig_{\circlearrowleft}\kappa_{a}]}{(\kappa_{m}J^{2}+\kappa_{a}\epsilon_{1}-\Delta_{a}\epsilon_{2})+i(\tilde{\Delta}_{m}J^{2}-2Jg_{\circlearrowright}g_{\circlearrowleft}+\Delta_{a}\epsilon_{1}+\kappa_{a}\epsilon_{2})},$ (A5)

### v4

@@ 6 Conclusion
We present a scheme for achieving nonreciprocal microwave -magnon and -phonon bipartite entanglements and photon-magnon-phonon tripartite entanglement in a CMM system based on the chiral cavity-magnon coupling. This is realized in a torus-shaped microwave cavity by placing the YIG sphere on a special line in the cavity to break the time-reversal symmetry of the two degenerate counter-propagating cavity modes. We analyze the effects of some major experimental imperfections, such as the coupling between the two degenerate cavity modes and the simultaneous couplings of the magnon mode to the two degenerate modes, and find that the nonreciprocity of the entanglements exhibits strong robustness towards these imperfections. Furthermore, we specifically provide a channel multiplexing quantum teleportation scheme based on the nonreciprocal entanglement. The work may find promising applications of the cavity magnomechanical systems in noise-tolerant quantum processing, channel multiplexing quantum teleportation, and chiral magnonic quantum networks.
Appendix A: Methods to verify magnon frequency combs
A strong driving field can generate magnon frequency combs [20]. To verify if this is present under the parameters we achieve entanglement, in the following we solve the differential equations to study the time evolution of the average of each mode. For completeness, here we include the effect of the backscattering induced coupling between the two degenerate counter-propagating cavity modes as in Sec. 4. The Hamiltonian of the system becomes
$\begin{split}&H^{\prime}/\hbar=\\ &\sum_{j=\circlearrowright,\circlearrowleft}\omega_{a}a_{j}^{\dagger}a_{j}+\omega_{m}m^{\dagger}m+\frac{\omega_{b}}{2}(q^{2}+p^{2})+g_{m}m^{\dagger}mq\\ &+\sum_{j=\circlearrowright,\circlearrowleft}g_{j}(a_{j}^{\dagger}m+a_{j}m^{\dagger})+J(a_{\circlearrowright}a_{\circlearrowleft}^{\dagger}+a_{\circlearrowright}^{\dagger}a_{\circlearrowleft})+H_{d}/\hbar,\end{split}$ (10)
where $J$ is the coupling strength between the two degenerate modes. When driving the CW mode of the cavity, the Hamiltonian (10) gives rise to the following differential equations for the classical averages of the modes
$\begin{split}\dot{\langle a_{\circlearrowright}\rangle}=&-i\Delta_{\circlearrowright}\langle a_{\circlearrowright}\rangle-\kappa_{a}\langle a_{\circlearrowright}\rangle-iJ\langle a_{\circlearrowleft}\rangle-ig_{\circlearrowright}\langle m\rangle+E,\\ \dot{\langle a_{\circlearrowleft}\rangle}=&-i\Delta_{\circlearrowleft}\langle a_{\circlearrowleft}\rangle-\kappa_{a}\langle a_{\circlearrowleft}\rangle-iJ\langle a_{\circlearrowright}\rangle-ig_{\circlearrowleft}\langle m\rangle,\\ \dot{\langle m\rangle}=&-i\Delta_{m}\langle m\rangle-\kappa_{m}\langle m\rangle-ig_{m}\langle m\rangle\langle q\rangle\\ &-ig_{\circlearrowright}\langle a_{\circlearrowright}\rangle-ig_{\circlearrowleft}\langle a_{\circlearrowleft}\rangle,\\ \dot{\langle q\rangle}=&\ \omega_{b}\langle p\rangle,\ \ \dot{\langle p\rangle}=-\omega_{b}\langle q\rangle-\gamma_{b}\langle p\rangle-g_{m}\langle m^{\dagger}\rangle\langle m\rangle.\\ \end{split}$ (11)
By solving the above equations and studying the time evolution of each mode, we can verify if the system enters a parameter regime where magnon frequency combs are generated.
Figure 8: The time evolution of the magnon amplitude $|\langle m(t)\rangle|$. Here we take $J=0.5\kappa_{m}$ and the other parameters are the same as used in Fig. 4(a).
Under the parameters of Fig. 4(a) and taking $J=0.5\kappa_{m}$, combining the corresponding initial conditions, we solve the time evolution of each mode, and Fig. 8 shows the result for the magnon mode amplitude $|\langle m(t)\rangle|$. It shows that the magnon amplitude evolves towards a stable value (similar to Fig. 2(a) of Ref. [20]), which implies that the drive power we used is below the threshold for producing magnon frequency combs. So, the presence of the frequency combs is excluded.
Appendix B: Linearized QLEs considering experimental imperfections
The Hamiltonian including the coupling between the two degenerate counter-propagating cavity modes is shown in Eq. (10). For the case where the microwave drive is applied to the CW circulating cavity mode, the linearized QLEs describing the quadrature fluctuations of the system can be written in the matrix form of
$\dot{u}(t)=A^{\prime}u(t)+n(t),$ (12)
$\begin{split}&A^{\prime}=\\ &\begin{pmatrix}-\kappa_{a}&\Delta_{\circlearrowright}&0&J&0&g_{\circlearrowright}&0&0\\ -\Delta_{\circlearrowright}&-\kappa_{a}&-J&0&-g_{\circlearrowright}&0&0&0\\ 0&J&-\kappa_{a}&\Delta_{\circlearrowleft}&0&g_{\circlearrowleft}&0&0\\ -J&0&-\Delta_{\circlearrowleft}&-\kappa_{a}&-g_{\circlearrowleft}&0&0&0\\ 0&g_{\circlearrowright}&0&g_{\circlearrowleft}&-\kappa_{m}&\tilde{\Delta}_{m}&\mathrm{Im}G_{m}&0\\ -g_{\circlearrowright}&0&-g_{\circlearrowleft}&0&-\tilde{\Delta}_{m}&-\kappa_{m}&-\mathrm{Re}G_{m}&0\\ 0&0&0&0&0&0&0&\omega_{b}\\ 0&0&0&0&-\mathrm{Re}G_{m}&-\mathrm{Im}G_{m}&-\omega_{b}&-\gamma_{b}\end{pmatrix},\end{split}$ (13)
$\begin{split}&\langle m\rangle=\\ &\frac{E[(g_{\circlearrowright}\Delta_{a}-g_{\circlearrowleft}J)-ig_{\circlearrowright}\kappa_{a}]}{(\kappa_{m}J^{2}+\kappa_{a}\epsilon_{1}-\Delta_{a}\epsilon_{2})+i(\tilde{\Delta}_{m}J^{2}-2Jg_{\circlearrowright}g_{\circlearrowleft}+\Delta_{a}\epsilon_{1}+\kappa_{a}\epsilon_{2})},\end{split}$ (14)
For the case where the CCW circulating cavity mode is driven, the drift matrix takes the same form as in Eq. (13), while the steady-state average $\langle m\rangle$ takes a slightly different form of
$\begin{split}&\langle m\rangle=\\ &\frac{E[(g_{\circlearrowleft}\Delta_{a}-g_{\circlearrowright}J)-ig_{\circlearrowleft}\kappa_{a}]}{(\kappa_{m}J^{2}+\kappa_{a}\epsilon_{1}-\Delta_{a}\epsilon_{2})+i(\tilde{\Delta}_{m}J^{2}-2Jg_{\circlearrowright}g_{\circlearrowleft}+\Delta_{a}\epsilon_{1}+\kappa_{a}\epsilon_{2})},\end{split}$ (15)
@@ Acknowledgements
We thank H. Jing for valuable feedback on the manuscript, and Z.-Q. Wang and R.-C. Shen for helpful discussions. This work was supported by National Key Research and Development Program of China (Grant No. 2022YFA1405200, 2024YFA1408900), National Natural Science Foundation of China (Grant No. 12474365, 92265202), and Zhejiang Provincial Natural Science Foundation of China (Grant No. LR25A050001).
Declaration of competing interest
The authors declare that they have no conflict of interest.